OpenStudy (anonymous):

Hadib is 8 years older then Cam. The sum of their ages is 24. step by step plz

OpenStudy (anonymous):

let age of cam be x yrs

OpenStudy (anonymous):

age of hadib = x+8 total age = 24 yrs x+ (x+8) = 24

OpenStudy (anonymous):

half of 24=12. 12+8= 20, and 12-8=4, so Cam is 4 and Hadib is 20